OpenStudy (anonymous):

-4x+y=6 -5x+y=21 Solving system of equations by substitutions?

OpenStudy (adamaero):

Solve for one variable and plug that into the other equation; substitute an x or y value to make one equation.

OpenStudy (anonymous):

-4x +y = 6 y = 6+4x -5x +y = 21 -5x + 6 +4x = 21 -x = 21-6 -x = 15 x = -15 y = 6+4x y = 6 + 4*(-15) y = 6 - 60 y = - 54

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

@adamaero that's what i try doing but didnt make sense and @sangya21 i dont think -54 is gonna plugin into the graph i have?

OpenStudy (anonymous):

@originaal_elizabeeth the method I showed above is the method of substitution As for graph, you can put x = 0,1, 2 and so on to find the value of y. and then plot that graph

OpenStudy (triciaal):

y = y 6 + 4 x = 21 + 5 x 6 - 21 = 5x - 4x x = -15 y = 6 + 4(-15) = 6 - 60 = - 54 check the original maybe you meant 2 and not 21 so that x = 4 with x = 4 y = 22

OpenStudy (triciaal):

by the way, what @sangya21 did is correct using substitution method

OpenStudy (anonymous):

this is the same thing im learning right now! ugh the struggle :/